OpenStudy (anonymous):

Solve on (0,2pi) interval. (sin x + 1)(2sin^2 x -3sin x - 2) option A. x=2pi , x = pi/2, x = pi/3 option B. x=pi, s=2pi/3, x = 5pi/3 option C. x= 3pi/2, x = 7pi/6, x = 11pi/6 option D. x = 2pi, x = pi/2, x = 5pi/4

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

Is the answer option B?

OpenStudy (anonymous):

@timo86m @ybarrap

OpenStudy (ybarrap):

@QueenBee232 , in problems like this where there is no equal sign or inequality, we can't solve. We can SIMPLIFY, but unless we know what this equals, we can solve. However, if the problem were, $$ (\sin x + 1)(2\sin^2 x -3\sin x - 2)=0 $$ Then that has a chance of having solutions because it equals something. One thing you CAN do is plot it on the given interval, but that is not the same thing as solving. Does that make sense? If the problem were given as I have stated above, the answer would NOT be B. You would be looking for solutions in the interval where each of the terms in parenthesis could be zero and B does not work.

OpenStudy (anonymous):

yeah in these problems they are looking for solutions for x with simplification

OpenStudy (anonymous):

so if its not B then I would have to say my second best guess is between A and D

OpenStudy (ybarrap):

Simplification and solving are two completely different things For example, $$ \sin x + 1=0\\ \implies \sin x = -1\\ \implies x=\cfrac{3\pi}{2}\\ $$ Then you need to solve $$ 2\sin^2 x -3\sin x - 2=0\\ $$ Here you can substitute y=sin x and use the quadratic formula to solve for sin x. Then determine the values of x that give you those 2 y values.

OpenStudy (ybarrap):

However, since x=3pi/2 is a solution if the problem is to solve for when your equation is zero, there is only one possible solution because only one has a x=3pi/2 as an answer.

OpenStudy (ybarrap):

Look at this problem - http://openstudy.com/study#/updates/53191234e4b0b3d1e04642e3 They ask the student to simplify. Look at the possible choices. None of the choices are solutions.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well I'm sorry to cut this short but I have to get going but I'll be back real soon

OpenStudy (anonymous):

and back